# Graph neural network outputs are almost surely asymptotically constant

### v1

@@ Abstract
Graph neural networks (GNNs) are the predominant architectures for a variety of learning tasks on graphs. We present a new angle on the expressive power of GNNs by studying how the predictions of a GNN probabilistic classifier evolve as we apply it on larger graphs drawn from some random graph model. We show that the output converges to a constant function, which upper-bounds what these classifiers can express uniformly. This convergence phenomenon applies to a very wide class of GNNs, including state of the art models, with aggregates including mean and the attention-based mechanism of graph transformers. Our results apply to a broad class of random graph models, including the (sparse) Erdős-Rényi model and the stochastic block model. We empirically validate these findings, observing that the convergence phenomenon already manifests itself on graphs of relatively modest size.
Convergence, uniform expressiveness, random graphs, graph neural networks, graph transformers, term language, machine learning, ICML
@@ 1 Introduction
Graphs are widely used for representing data in a broad range of domains. Graph neural networks (GNNs) (Scarselli et al., 2009; Gori et al., 2005) have become prominent for graph ML for many tasks with successful applications in life sciences (Shlomi et al., 2021; Duvenaud et al., 2015; Kearnes et al., 2016; Zitnik et al., 2018) among others. Most GNN architectures are instances of message passing neural networks (MPNNs) (Gilmer et al., 2017). Recently, there has been a surge in the use of graph transformer architectures (Ying et al., 2021; Rampášek et al., 2022; Ma et al., 2023) aiming to transfer the success of vanilla transformers from language processing  (Vaswani et al., 2017) to graphs.
The empirical success of graph representation learning architectures motivated a large body of work investigating their expressive power. In this paper, we build on a rich line work (Keriven et al., 2021, 2020; Maskey et al., 2022; Levie, 2023; Adam-Day et al., 2023) on the convergence properties of GNNs. We consider graph learning architectures acting as probabilistic classifiers (e.g., using softmax) and ask how the predictions of such classifiers evolve as we apply them on larger graphs drawn from a random graph model.
We show that the output of many common probabilistic classifiers eventually become independent of their inputs as graph size increases. In other words, each model eventually outputs the same prediction probabilities on all graphs. We show this convergence phenomena to be extremely robust in the following two dimensions.
Firstly, our results apply to a broad class of graph learning architectures with different aggregate and update functions, encompassing GNNs such as graph attention networks (Veličković et al., 2018), as well as popular (graph) transformers, such as the General, Powerful, Scalable Graph Transformer with random walk encodings (Rampášek et al., 2022). The flexibility of our results stems from our use of a term language, which allows combining graph operations arbitrarily.
@@
Secondly, all presented results apply to a broad class of random graph models, including Erdős-Rényi models of various sparsity levels, and the stochastic block model. The sparse models are more realistic than their dense counterparts which makes it typically harder to obtain results for them. This is also reflected in our study, as the results for sparse and dense models require very different proofs.
The convergence phenomenon provides an upper bound on the uniform expressiveness of any architecture which can be expressed with our term language. The immediate practical consequence is that the only probabilistic classifiers which can be uniformly expressed by these architectures are those which are asymptotically constant.
The technical machinery introduced in this paper could be of independent interest, since we introduce an aggregate term language, which has attractive closure properties (Section 4), and provide an “almost sure optimization” result, saying that terms in the language can be simplified to Lipschitz functions for most inputs (Section 5). We validate these results empirically, showing the convergence of these probabilistic classifiers in practice (Section 6). We observe rapid convergence for all graph architectures and across all graph distributions considered. Interestingly, we note some distinctions between the convergence of the sparse and non-sparse Erdős-Rényi model, which we can relate to the proof strategies for our convergence laws.
@@ 2 Related work
Uniform expressiveness. The expressive power of MPNNs is studied from different angles, including their power in terms of graph distinguishability (Xu et al., 2019; Morris et al., 2019), counting subgraphs (Chen et al., 2020), expressing logical (node) classifiers (Barcelo et al., 2020), or detecting graph (bi)connectivity (Zhang et al., 2023). The seminal results of Xu et al. (2019); Morris et al. (2019) show that MPNNs are upper bounded by the 1-dimensional Weisfeiler Leman graph isomorphism test (1-WL) in terms of graph distinguishability. One subtle aspect regarding the WL-style expressiveness results is that they are inherently non-uniform, i.e., the model construction is dependent on the graph size. One notable exception is the work by Barcelo et al. (2020) who present a uniform expressiveness result. There are also recent studies that focus on uniform expressiveness (Rosenbluth et al., 2023; Adam-Day et al., 2023). In particular, Adam-Day et al. (2023) investigate the uniform expressive power of GNNs with randomized node features, which are known to be more expressive in the non-uniform setting (Abboud et al., 2021; Sato et al., 2021). They show that for classical Erdős-Rényi graphs, GNN binary classifiers display a zero-one law, assuming certain restrictions on GNN weights and the random graph model. We focus on probabilistic classifiers, where their results do not apply, and provide results for a wider class of random graph models, subsuming state of the art architectures such as graph transformers.
Convergence laws for languages. Our work situates GNNs within a rich term language built up from graph and node primitives via real-valued functions and aggregates. Thus it relates to convergence laws for logic-based languages on random structures, dating back to the zero-one law of Fagin (1976) for Erdős-Rényi random graphs with fixed probability. This line includes results for uniform distributions on restricted classes of structures (Lynch, 1993), and sparse variants of the Erdős-Rényi random graph model (Shelah & Spencer, 1988; Larrauri et al., 2022). While the bulk of these results are for first-order logic, there are also strong concentration results for subsets of infinitary logic (Kolaitis & Vardi, 1992). We know of no prior convergence laws for languages with aggregates; the only work on numerical term languages is by Grädel et al. (2022), which deals with a variant of first-order logic in general semi-rings.
Other notions of convergence on random graphs. The works of Cordonnier et al. (2023); Keriven et al. (2021, 2020); Maskey et al. (2022); Levie (2023) consider convergence to continuous analogues of GNNs, often working within metrics on a function space. Our approach is fundamentally different in that we can use the standard notion of asymptotic convergence in Euclidean space, comparable to traditional language-based convergence results outlined above, such as those by Fagin (1976) and Lynch (1992). Moreover, for our weighted mean-based term language, we obtain both broad and simple criteria for convergence, compared to the technically intricate criteria given by Cordonnier et al. (2023).
@@ 3.1 Featured random graphs and convergence
Random graphs. We consider simple, undirected graphs $G=(V_{G},E_{G},H_{G})$ where each node is associated with a vector of node features given by $H_{G}:V_{G}\to{\mathbb{R}}^{d}$. We refer to this as a featured graph. We commonly refer to a tuple of nodes denoted as $\bar{v}$. We are interested in random graph models, specifying for each number $n$ a distribution $\mu_{n}$ on graphs with $n$ nodes, along with random graph feature models, where we have a distribution $\mu_{n}$ on featured graphs with $n$ nodes. Given a random graph model and a distribution $\nu$ over $\mathbb{R}^{d}$, we get a random graph feature model by letting the node features be chosen independently of the graph structure via $\nu$.
Erdős-Rényi and Stochastic Block Model. The most basic random graph model we deal with is the Erdős-Rényi $\mathrm{ER}(n,p(n))$, where an edge is included in the graph with $n$ nodes with probability $p(n)$. The classical case is when $p(n)$ is independent of the graph size $n$, which we refer as the dense ER distribution. We also consider the stochastic block model $\mathrm{SBM}(n_{1},\ldots,n_{M},P)$, which contains $m$ communities of sizes $n_{1},\ldots,n_{M}$ and an edge probability matrix between communities ${\bm{P}}\in{\mathbb{R}}^{M\times M}$. A community $i$ is sampled from the Erdős-Rényi distribution $\mathrm{ER}(n,p(n)={\bm{P}}_{i,i})$ and an edge to a node in another community $j$ is included with probability ${\bm{P}}_{i,j}$.
Almost sure convergence. Given any function $F$ from featured graphs to real vectors and a random featured graph model $\bar{\mu}=\mu_{n}:n\in\mathbb{N}$, we say $F$ converges asymptotically almost surely (converges a.a.s.) to a vector ${\bm{z}}$ with respect to $\bar{\mu}$ if for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$, with probability at least $1-\theta$ when drawing featured graphs $G$ from $\mu_{n}$, we have that $\lVert F(G)-{\bm{b}}\rVert<\epsilon$. We extend this terminology to a family of functions $\mathcal{F}$ as above and a random featured graph model $\bar{\mu}$, saying $\mathcal{F}$ is $\bar{\mu}$-convergent if all functions in $F$ converge a.a.s. w.r.t. $\bar{\mu}$.
@@ 3.2 Graph neural networks and graph transformers
We first briefly introduce framework of message passing neural networks (MPNNs)  (Gilmer et al., 2020), which includes the vast majority of graph neural networks and then introduce (graph) transformers.
@@
We denote by $d^{(\ell)}$ the dimension of the node embeddings at iteration (layer) $\ell$. The final node representations ${\bm{h}}_{v}^{\left(L\right)}$ of each node $v$ can then be used for node-level predictions. For graph-level predictions, the final node embeddings are pooled to form a graph embedding vector ${\bm{z}}_{G}^{\left(L\right)}$ to predict properties of entire graphs, which can be, e.g., averaging, summation, or element-wise maximum.
GATs. One MPNN instance we examine is graph attention networks (GATs) (Veličković et al., 2018), where each node is updated with a weighted average of its neighbors’ representations, letting ${\bm{h}}_{v}^{\left(\ell+1\right)}$ be:
@@
where $\mathrm{score}({\bm{h}}_{v}^{(\ell)},{\bm{h}}_{u}^{(\ell)})$ is given by:
$\text{LeakyReLU}\left({{\bm{a}}^{(\ell)}}^{\top}[{\bm{W}}^{(\ell)}{\bm{h}}_{v}^{(\ell)}||{\bm{W}}^{(\ell)}{\bm{h}}_{u}^{(\ell)}]\right),$
and ${\bm{a}}^{(\ell)}\in\mathbb{R}^{2^{\left(\ell\right)}},{\bm{W}}^{(\ell)}\in\mathbb{R}^{d^{\left(\ell+1\right)}\times d^{\left(\ell\right)}}$ are learnable parameters, and ||||| | denotes vector concatenation.
@@
$\small\textsc{Att}(v)=\sum_{u\in V_{G}}\frac{{\exp\left(\mathrm{scale}\left({\bm{h}}_{v}^{(\ell)},{\bm{h}}_{u}^{(\ell)}\right)\right)}}{\sum\limits_{w\in V_{G}}\exp\left(\mathrm{scale}\left({\bm{h}}_{v}^{(\ell)},{\bm{h}}_{w}^{(\ell)}\right)\right)}{\bm{W}}^{(\ell)}_{V}{\bm{h}}_{u}^{(\ell)},$
where $\mathrm{scale}$ is the usual scaled dot-product attention:
$\mathrm{scale}\left({\bm{h}}_{v}^{(\ell)},{\bm{h}}_{u}^{(\ell)}\right)=\frac{\left({{\bm{W}}^{(\ell)}_{Q}{\bm{h}}_{v}^{(t)}}\right)^{\top}\left({{\bm{W}}^{(\ell)}_{K}{\bm{h}}_{u}^{(t)}}\right)}{\sqrt{d_{k}}},$
and ${\bm{W}}_{Q}^{\left(\ell\right)}$, ${\bm{W}}_{K}^{\left(\ell\right)}$, and ${\bm{W}}_{V}^{\left(\ell\right)}$ are layer-specific linear maps respectively acting on query, key, and value nodes.
The vanilla transformer architecture ignores the graph structure. Graph transformer architectures (Ying et al., 2021; Rampášek et al., 2022; Ma et al., 2023) address this by explicitly encoding graph inductive biases, most typically in the form of positional encodings (PEs). In their simplest form, these encodings are additional features ${\bm{p}}_{v}$ for every node $v$ that encode a node property (e.g., node degree) which are concatenated to the node features ${\bm{h}}_{v}$. The most common positional encodings are Laplacian positional encodings and random walk positional encodings (RW) (Dwivedi et al., 2021). For each node, encoding is given by:
@@
The GPS architecture (Rampášek et al., 2022) is a representative graph transformer, which applies a parallel computation in every layer: a transformer layer (with or without PEs) and an MPNN layer are applied in parallel and their outputs are summed to yield the node representations. By including a standard MPNN in this way, a GPS layer can take advantage of the graph topology even when there is no positional encoding. In the context of this paper, we write GPS to refer to a GPS architecture that uses an MPNN with mean aggregation, and GPS+RW if the architecture additionally uses a random-walk PE.
Probabilistic classifiers. All of the models defined above can be used for probabilistic graph-level classification. We only need to ensure that the final layer is a softmax or sigmoid applied to pooled representations.
@@ 4 Model architectures via term languages
4.1 Term languages on graphs
We demonstrate the robustness and generality of the convergence phenomenon by defining a term language consisting of compositions of operators on graphs. Terms are formal sequences of symbols which when interpreted in a given graph yield a real-valued function on the graph nodes. Our language subsumes a huge variety of GNNs.
@@ 4.1 Term languages on graphs
  • The basic terms are of the form $\mathrm{H}(x)$, which returns the features the node corresponding to variable $x$, and constants ${\bm{c}}$.
  • Let $h\colon\mathbb{R}^{d}\to(0,\infty)^{d}$ be a function which is Lipschitz continuous on every compact domain. Given terms $\tau$ and $\pi$, the local $h$-weighted mean for node variable $x$ is:
@@
    The global $h$-weighted mean is the term:
@@
  • Terms are closed under applying a function symbol for each Lipschitz continuous $F\colon\mathbb{R}^{d\times k}\to\mathbb{R}^{d}$.
Thus the weighted mean operator takes an average of the values returned by $\pi$, using $\tau$ to perform a weighting and normalizing the values of $\pi$ using $h$ to ensure that we are not dividing by zero.
@@
  • $\llbracket{\bm{c}}(\bar{u})\rrbracket_{G}={\bm{c}}$ for any constant ${\bm{c}}$.
  • $\llbracket\mathrm{H}(x_{i})(\bar{u})\rrbracket_{G}=H_{G}(u_{i})$, the $i$th node’s features.
  • $\llbracket F(\tau_{1},\ldots,\tau_{k})(\bar{u})\rrbracket_{G}=F(\llbracket\tau_{1}(\bar{u})\rrbracket_{G},\ldots,\llbracket\tau_{k}(\bar{u})\rrbracket_{G})$ for any function symbol $F$.
  • Define $\left\llbracket\sum_{y\in\mathcal{N}(x_{i})}\tau\star h(\pi)(\bar{u})\right\rrbracket_{G}$ as:
    $\frac{\sum\limits_{v\in\mathcal{N}(u_{i})}\llbracket\tau(\bar{u},v)\rrbracket_{G}h(\llbracket\pi(\bar{u},v)\rrbracket_{G})}{\sum\limits_{v\in\mathcal{N}(u_{i})}h(\llbracket\pi(\bar{u},v)\rrbracket_{G})}$
    when $\operatorname{\mathcal{N}}(u_{i})$ is non-empty and ${\bf 0}$ otherwise.
  • The semantics of global weighted mean is defined analogously.
A closed term has all node variables bound by the Wmean operator: so the implicit input is just a featured graph.

### v2

# Almost Surely Asymptotically Constant Graph Neural Networks

@@ Abstract
We present a new angle on the expressive power of graph neural networks (GNNs) by studying how the predictions of a GNN probabilistic classifier evolve as we apply it on larger graphs drawn from some random graph model. We show that the output converges to a constant function, which upper-bounds what these classifiers can uniformly express. This strong convergence phenomenon applies to a very wide class of GNNs, including state of the art models, with aggregates including mean and the attention-based mechanism of graph transformers. Our results apply to a broad class of random graph models, including sparse and dense variants of the Erdős–Rényi model, the stochastic block model, and the Barabási-Albert model. We empirically validate these findings, observing that the convergence phenomenon appears not only on random graphs but also on some real-world graphs.
@@ 1 Introduction
Figure 1: The output of the considered GNNs eventually become constant as the graph sizes increase.
Graph neural networks (GNNs) [47, 21] have become prominent for graph ML with applications in domains such as life sciences [50, 11, 25, 58]. Their empirical success motivated work investigating their theoretical properties, pertaining to their expressive power [55, 42, 4, 8, 1, 46, 23, 19, 57], generalization capabilities [18, 33, 40, 48, 43], and convergence properties [28, 27, 40, 32, 2].
We consider GNNs acting as probabilistic graph classifiers and ask: how do the predictions of GNNs acting as probabilistic graph classifiers evolve as we apply them on larger graphs drawn from a random graph model?
Our study gives a surprising answer to this question: the output of many probabilistic classifiers eventually become independent of their inputs – each model eventually outputs the same prediction probabilities on all graphs – as graph sizes increase (Figure 1). This “almost sure convergence” to a constant distribution is much stronger than convergence to some limit object [28, 27, 32]. The immediate consequence of this strong convergence phenomenon is to upper bound the uniform expressiveness of the considered model architectures: these architectures can uniformly express only the probabilistic classifiers that are almost surely asymptotically constant.
Scope of the result. The core approach in graph ML is based on iteratively updating node representations of an input graph by an aggregate of messages flowing from the node’s neighbours [20]. This approach can be extended to global aggregates [5]. Our main result holds for all architectures that use weighted mean as an aggregation function and it is extremely robust in the following two dimensions:
  1. 1.
    Model architectures: Our result is very general and abstracts away from low-level architectural design choices. To achieve this, we introduce an aggregate term language using weighted mean aggregation and provide an “almost sure optimization” result for this language: our result states that every term in the language can be simplified to Lipschitz functions for most inputs. Thus, any architecture that can be expressed in this language follows the same convergence law. This includes graph attention networks (GATs) [52], as well as popular (graph) transformers, such as the General, Powerful, Scalable Graph Transformer (GPS) with random walk encodings [44]. The term language can seemlessly capture common design choices, such as skip and jumping knowledge connections [54], or global aggregation schemes [5].
  2. 2.
@@
    Random graph models: All results apply to a wide class of random graph models, including Erdős-Rényi models of various sparsity levels, the Barabási-Albert preferential attachment model, and the stochastic block model. The sparse models are more realistic than their dense counterparts which makes it typically harder to obtain results for them. This is also reflected in our study, as the results for sparse and dense models require very different proofs.
Contributions. The key contributions of this paper are as follows:
  • We introduce a flexible aggregate term language with attractive closure properties (Section 4) and prove an “almost sure convergence” result for this language relative to a wide class of random graph models (Section 5). This result is of independent interest since it pushes the envelope for convergence results from classical logical languages [15, 36] to include aggregation.
  • We show that a diverse class of architectures acting as probabilistic graph classifiers can be expressed in the term language (Section 4) and derive “almost sure convergence” as corollaries to our main result (Corollary 5.2). The results are robust to many practical architectural design choices and even hold for architectures using mixtures of layers from different architectures.
  • We validate these results empirically, showing the convergence of these probabilistic graph classifiers in practice (Section 6) on graphs drawn from the random models studied. In addition, we probe the real-world import of our results by testing for convergence on a dataset with varying size dataset splits. Across all experiments we observe rapid convergence to a constant distribution. Interestingly, we note some distinctions between the convergence of the sparse and non-sparse Erdős-Rényi model, which we can relate to the proof strategies for our convergence laws.
@@ 2 Related work
Uniform expressiveness. The expressive power of MPNNs is studied from different angles, including their power in terms of graph distinguishability [55, 42]. The seminal results of Xu et al. [55], Morris et al. [42] show that MPNNs are upper bounded by the 1-dimensional Weisfeiler Leman graph isomorphism test (1-WL) in terms of graph distinguishability. WL-style expressiveness results are inherently non-uniform, i.e., the model construction is dependent on the graph size. There are also recent studies that focus on uniform expressiveness [4, 45, 2]. In particular, Adam-Day et al. [2] investigate the uniform expressive power of GNNs with randomized node features, which are known to be more expressive in the non-uniform setting [1, 46]. They show that for classical Erdős-Rényi graphs, GNN binary classifiers display a zero-one law, assuming certain restrictions on GNN weights and the random graph model. We focus on probabilistic classifiers, where their results do not apply, while dealing with a wider class of random graph models, subsuming popular architectures such as graph transformers.
Convergence laws for languages. Our work situates GNNs within a rich term language built up from graph and node primitives via real-valued functions and aggregates. Thus it relates to convergence laws for logic-based languages on random structures, dating back to the zero-one law of Fagin [15], including [35, 49, 31]. We are not aware of any prior convergence laws for languages with aggregates; the only work on numerical term languages is by Grädel et al. [22], which deals with a variant of first-order logic in general semi-rings.
Other notions of convergence on random graphs. The works of Cordonnier et al. [9], Keriven et al. [28, 27], Maskey et al. [40], Levie [32] consider convergence to continuous analogues of GNNs, often working within metrics on a function space. The results often focus on dense random graph models, such as graphons [34]. Our approach is fundamentally different in that we can use the standard notion of asymptotic convergence in Euclidean space, comparable to traditional language-based convergence results outlined above, such as those by Fagin [15] and Lynch [36]. The key point is that a.a.s. constancy is a very strong notion of convergence and it does not follow from convergence in the senses above. In fact, obtaining such a strong convergence result depends heavily on the details of the term language, as well as the parameters that control the random graph: see Section 7 for further discussion of the line between a.a.s. convergence and divergence. Our study gives particular emphasis to sparse random graph models, like Barabási-Albert, which are closer to graphs arising in practice.
@@ 3.1 Featured random graphs and convergence
Random graphs. We consider simple, undirected graphs $G=(V_{G},E_{G},H_{G})$ where each node is associated with a vector of node features given by $H_{G}:V_{G}\to{\mathbb{R}}^{d}$. We refer to this as a featured graph. We are interested in random graph models, specifying for each number $n$ a distribution $\mu_{n}$ on graphs with $n$ nodes, along with random graph feature models, where we have a distribution $\mu_{n}$ on featured graphs with $n$ nodes. Given a random graph model and a distribution $\nu$ over $\mathbb{R}^{d}$, we get a random graph feature model by letting the node features be chosen independently of the graph structure via $\nu$.
Erdős-Rényi and the stochastic block model. The most basic random graph model we deal with is the Erdős-Rényi distribution $\mathrm{ER}(n,p(n))$, where an edge is included in the graph with $n$ nodes with probability $p(n)$ [14]. The classical case is when $p(n)$ is independent of the graph size $n$, which we refer as the dense ER distribution. We also consider the stochastic block model $\mathrm{SBM}(n_{1},\ldots,n_{M},P)$, which contains $m$ communities of sizes $n_{1},\ldots,n_{M}$ and an edge probability matrix between communities ${\bm{P}}\in{\mathbb{R}}^{M\times M}$. A community $i$ is sampled from the Erdős-Rényi distribution $\mathrm{ER}(n,p(n)={\bm{P}}_{i,i})$ and an edge to a node in another community $j$ is included with probability ${\bm{P}}_{i,j}$.
The Barabási-Albert preferential attachment model. Many graphs encountered in the real world obey a power law, in which a few vertices are far more connected than the rest [7]. The Barabási-Albert distribution was developed to model this phenomenon [3]. It is parametrised by a single integer $m$, and the $n$-vertex graph $\mathrm{BA}(n,m)$ is generated sequentially, beginning with a fully connected $m$-vertex graph. Nodes are added one at a time and get connected via $m$ new edges to previous nodes, where the probability of attaching to a node is proportional to its degree.
Almost sure convergence. Given any function $F$ from featured graphs to real vectors and a random featured graph model $(\mu_{n})_{n\in\mathbb{N}}$, we say $F$ converges asymptotically almost surely (converges a.a.s.) to a vector ${\bm{z}}$ with respect to $\bar{\mu}$ if for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$, with probability at least $1-\theta$ when drawing featured graphs $G$ from $\mu_{n}$, we have that $\lVert F(G)-{\bm{z}}\rVert<\epsilon$.
@@ 3.2 Graph neural networks and graph transformers
We first briefly introduce message passing neural networks (MPNNs)  [20], which include the vast majority of graph neural networks, as well as (graph) transformers.
@@
GATs. One class of MPNNs are graph attention networks (GATs) [52], where each node is updated with a weighted average of its neighbours’ representations, letting ${\bm{h}}_{v}^{\left(\ell+1\right)}$ be:
@@
where $\mathrm{score}$ is a certain learnable Lipschitz function.
@@
$\small\textsc{Att}(v)=\sum_{u\in V_{G}}\frac{{\exp\left(\mathrm{scale}\left({\bm{h}}_{v}^{(\ell)},{\bm{h}}_{u}^{(\ell)}\right)\right)}}{\sum\limits_{w\in V_{G}}\exp\left(\mathrm{scale}\left({\bm{h}}_{v}^{(\ell)},{\bm{h}}_{w}^{(\ell)}\right)\right)}{\bm{W}}^{(\ell)}_{V}{\bm{h}}_{u}^{(\ell)}$
where $\mathrm{scale}$ is another learnable Lipschitz function (the scaled dot-product).
The vanilla transformer architecture ignores the graph structure. Graph transformer architectures [56, 44, 37] address this by explicitly encoding graph inductive biases, most typically in the form of positional encodings (PEs). In their simplest form, these encodings are additional features ${\bm{p}}_{v}$ for every node $v$ that encode a node property (e.g., node degree) which are concatenated to the node features ${\bm{h}}_{v}$. The random walk positional encoding (RW) [12] of each node $v$ is given by:
@@
The GPS architecture [44] is a representative graph transformer, which applies a parallel computation in every layer: a transformer layer (with or without PEs) and an MPNN layer are applied in parallel and their outputs are summed to yield the node representations. By including a standard MPNN in this way, a GPS layer can take advantage of the graph topology even when there is no positional encoding. In the context of this paper, we write GPS to refer to a GPS architecture that uses an MPNN with mean aggregation, and GPS+RW if the architecture additionally uses a random-walk PE.
Probabilistic classifiers. All of these models can be used as probabilistic graph classifiers. We only need to ensure that the final layer is a softmax or sigmoid applied to pooled representations.
@@ 4 Model architectures via term languages
We demonstrate the robustness and generality of the convergence phenomenon by defining a term language consisting of compositions of operators on graphs. Terms are formal sequences of symbols which when interpreted in a given graph yield a real-valued function on the graph nodes.
@@
  1. The basic terms are of the form $\mathrm{H}(x)$, representing the features of the node $x$, and constants ${\bm{c}}$.
  2. Let $h\colon\mathbb{R}^{d}\to(0,\infty)^{d}$ be a function which is Lipschitz continuous on every compact domain. Given terms $\tau$ and $\pi$, the local $h$-weighted mean for node variable $x$ is:
@@
    The interpretation of $\star$ will be defined below. The global $h$-weighted mean is the term:
@@
  3. Terms are closed under applying a function symbol for each Lipschitz continuous $F\colon\mathbb{R}^{d\times k}\to\mathbb{R}^{d}$.
The weighted mean operator takes an average of the values returned by $\pi$, using $\tau$ to perform a weighting and normalizing the values of $\pi$ using $h$ to ensure that we are not dividing by zero.
@@
  1. $\llbracket{\bm{c}}(\bar{u})\rrbracket_{G}={\bm{c}}$ for any constant ${\bm{c}}$.
  2. $\llbracket\mathrm{H}(x_{i})(\bar{u})\rrbracket_{G}=H_{G}(u_{i})$ for the $i$th node’s features.
  3. $\llbracket F(\tau_{1},\ldots,\tau_{k})(\bar{u})\rrbracket_{G}=F(\llbracket\tau_{1}(\bar{u})\rrbracket_{G},\ldots,\llbracket\tau_{k}(\bar{u})\rrbracket_{G})$ for any function symbol $F$.
  4. For any term composed using $\star$:
    $\left\llbracket\sum_{y\in\mathcal{N}(x_{i})}\tau\star h(\pi)(\bar{u})\right\rrbracket_{G}=\left\{\begin{array}[]{cc}\nicefrac{{\sum\limits_{v\in\mathcal{N}(u_{i})}\llbracket\tau(\bar{u},v)\rrbracket_{G}h(\llbracket\pi(\bar{u},v)\rrbracket_{G})}}{{\sum\limits_{v\in\mathcal{N}(u_{i})}h(\llbracket\pi(\bar{u},v)\rrbracket_{G})}}&\mbox{if $\operatorname{\mathcal{N}}(u_{i})\neq\emptyset$};\\ {\bf 0}&\mbox{if $\operatorname{\mathcal{N}}(u_{i})=\emptyset$}.\end{array}\right.$
The semantics of global weighted mean is defined analogously and omitted for brevity.
A closed term has all node variables bound by a weighted mean operator: so the implicit input is just a featured graph.